# Zero-resource Hallucination Detection for Text Generation via Graph-based Contextual Knowledge Triples Modeling

### v3

@@ Abstract
LLMs obtain remarkable performance but suffer from hallucinations. Most research on detecting hallucination focuses on the questions with short and concrete correct answers that are easy to check the faithfulness. Hallucination detections for text generation with open-ended answers are more challenging. Some researchers use external knowledge to detect hallucinations in generated texts, but external resources for specific scenarios are hard to access. Recent studies on detecting hallucinations in long text without external resources conduct consistency comparison among multiple sampled outputs. To handle long texts, researchers split long texts into multiple facts and individually compare the consistency of each pairs of facts. However, these methods (1) hardly achieve alignment among multiple facts; (2) overlook dependencies between multiple contextual facts. In this paper, we propose a graph-based context-aware (GCA) hallucination detection for text generations, which aligns knowledge facts and considers the dependencies between contextual knowledge triples in consistency comparison. Particularly, to align multiple facts, we conduct a triple-oriented response segmentation to extract multiple knowledge triples. To model dependencies among contextual knowledge triple (facts), we construct contextual triple into a graph and enhance triples’ interactions via message passing and aggregating via RGCN. To avoid the omission of knowledge triples in long text, we conduct a LLM-based reverse verification via reconstructing the knowledge triples. Experiments show that our model enhances hallucination detection and excels all baselines.  [^1] The code is at: github.com/GCA-hallucinationdetection/GCA
Introduction
@@
Recent research showed that large language models (LLMs) achieved state-of-the-art performance in various NLP tasks (Qin et al. 2023; Fang et al. 2024). However, these models often suffer from hallucinations: generate incorrect or fabricated content in a factual way, which undermines models’ credibility (Ji et al. 2023) and limits LLMs’ application in fields requiring factual accuracy (Huang et al. 2023). Detecting hallucination in model responses is crucial for LLMs’ boom.
@@
Existing hallucination detection studies primarily focus on tasks (e.g. question answering (QA) (Zhang et al. 2023) and arithmetic calculation (Xue et al. 2023)), where the questions have short and concrete correct answers. In these tasks, the consistency among their concrete answers can be easily checked for hallucination (Jiang et al. 2021). The research on detecting hallucination on generated long text is more challenging because (1) generating text is open-ended and seldom has concrete answers; and (2) the long text contains multiple facts and the consistency among multiple facts is hard to verify. Some studies use external knowledge to verify the factuality of outputted responses (Mishra et al. 2024; Li et al. 2024a; Sadat et al. 2023), but high-quality external resources for specific tasks are expensive and hard to access (Luo, Xiao, and Ma 2023). Therefore, it is valuable to investigate hallucination detection for long text generation without external resources (i.e. zero-resource setting). Hence, in this paper, we aim to explore hallucination detection for long text generation with black-box models (powerful LLMs, like GPT-4) in a zero-resource setting.
@@
Currently, the studies on black-box zero-resource hallucination detection for text generation can be divided into two categories: (1) Self-checking (Friel and Sanyal 2023; Liu et al. 2023) designs prompt texts using chain-of-thought (CoT) to check whether models’ responses contains hallucinations by the LLMs’ own capabilities. Though that can be easily implemented and simply applied in various scenarios, it relies on the model’s own ability and thus probably leads to overconfidence: The model tends to believe that its own responses are correct, leading to missed detection. (2) Consistency comparison (Zhang et al. 2023; Ma, Dai, and Sui 2024) samples multiple responses to check whether the sampled responses are highly inconsistent, which indicates confusion (hallucination) about the facts in responses (Farquhar et al. 2024). The method is effective when the responses are very short with few concrete claims, which are easy to compare consistency in the sampled responses. In addition, to detect hallucinations in long response texts, multiple responses have various wording or lexical representations about a same knowledge fact or term (Donald Trump vs Donald John Trump), which results in the difficulty of comparing facts claimed in multiple responses.
@@
To address those issues in hallucination detection for long texts, researchers proposed divide-and-conquer (Zhang et al. 2024) based on consistency comparison. It has three steps: (1) sample multiple additional responses appending to the original response to the user’s query; (2) divide each response into multiple facts; (3) compare the consistency of facts in the original response with those in the sampled responses, where highly inconsistent facts indicates hallucinations. However, this method ignores the omission issue: the facts, which are in the original response but not in the sampled responses, may be falsely identified as hallucinations due to the lack of comparison. To overcome that, some researchers proposed reverse verification (Yang, Sun, and Wan 2023; Cao, Yang, and Zhao 2023) based on the divide-and-conquer approach. For each fact in the original response, they prompt LLMs to create a new question answered by the claim, and compare the model’s response consistency regarding the knowledge fact. The method can avoid the omission issue because each fact corresponds to an accurately matched counterpart for comparison. However, a long text may contain multiple knowledge facts $\{A_{1},...A_{N}\}$, where $A_{i}$ is the $i$-th fact within LLMs’ generation. The above method compares the consistency of each single fact sampled from different responses individually (e.g. verify the truthfulness of $A_{i}$ by comparing with $B_{i}$, where $B_{i}$ is $A_{i}$’s corresponding fact from a sampled response). However, this approach ignores the dependencies among each fact (e.g. $A_{i-1}$ and $A_{i}$ is in a response and may have be dependency). The lack of dependency information may result in suboptimal performance while detecting the truthfulness of individual facts. Therefore, considering the relationship (inherent in the context around the facts) between a single fact and its surrounding facts is promising to enhance hallucination detection, these relationships are inherent in the context around the facts.
@@
In this paper, we propose a graph-based context-aware (GCA) hallucination detection method for long text generation, which abstracts to align facts for a better consistency comparison and considers the dependencies between contextual knowledge triples. Specifically, to better align facts from responses for consistency comparison, we conduct a triple-oriented response segmentation to extract multiple knowledge triples for each response. Triples are well-structured and thus easy for comparison; in addition, it is easy to align facts among multiple responses since triples carry only knowledge facts and get rid of the wording and lexical representations in the response(Wang et al. 2023). (Sec. 3.2). To consider dependencies between a triple and its contextual triples, we construct a graph over a triple and its contextual triples, and then encourage triples’ interactions via message passing and aggregating the features of neighboring nodes via RGCN(Wang et al. 2020) (Sec. 3.3). It encourages the interaction between knowledge facts as we observe that a certain number of facts has dependencies in a response (See Sec. 4.5). To avoid the omission issue in (Sec. 3.3), we propose a LLM-based reverse verification method with three reconstruction tasks to reconstruct the knowledge triples. (Sec. 3.4). Experiments show that our method effectively improves hallucination detection accuracy for long text responses generated by black-box models under zero-resource conditions.
@@
Our contributions are: (1) we propose to consider the dependencies between contextual knowledge triples hallucination detection method for long text generation. (2) We propose a graph-based context-aware hallucination detection via consistency comparison with RGCN. (3) We additionally proposed three reversion verification tasks to helps hallucination detection by reconstructing triples. (4) Experiments show that our method outperforms all baselines.
Related Work
White-box Hallucination Detection
@@
These methods identify hallucinations by analyzing the internal states of the model (Yuksekgonul et al. 2023), which is classified mainly into two types: (1) Output logit based method: The model’s output logits reflect the confidence of the model’s predictions. (Jiang et al. 2024). (Varshney et al. 2023) calculate the logits for concepts in response and take the minimal probabilities to model the uncertainty. (Verma et al. 2023) explore the integration of Epistemic Neural Networks (ENNs) with LLMs to improve the model’s uncertainty estimates. (Luo, Xiao, and Ma 2023) proposed to adjust the model’s output logits by adding a linear layer to better align with the actual correctness likelihood. (2) Hidden layer activations based method: Hidden layer activations encapsulate the model’s internal representation of statement truthfulness (Fadeeva et al. 2024). (Azaria and Mitchell 2023) train a classifier using LLM’s hidden layer activations to assess the truthfulness of statements. (Snyder, Moisescu, and Zafar 2023) use output values from artifacts associated with model generation as input features to train a classifier that identifies hallucinations. (Zhu et al. 2024) uses probabilistic models to analyze internal state transitions in the LLM during generation to detect hallucinations.
Black-box Hallucination Detection using External Resources
@@
These methods aim to verify the authenticity of model-generated content by leveraging external knowledge (Manakul, Liusie, and Gales 2023; Nahar et al. 2024). Depending on the source of external knowledge, it can be categorized into the following two types. (1) RAG-based method: Retrieval-augmented generation (RAG) is a technique that enhances text generation by retrieving relevant information from external sources (Sadat et al. 2023). (Roychowdhury et al. 2024) propose a multi-LLM system with the capability to retrieve external knowledge bases and perform real-time content authentication. (Ding et al. 2024) retrieve relevant evidence to help LLMs correct potential hallucinations in responses. (Kang, Ni, and Yao 2023) propose a method for the real-time retrieval of Web search engines that can verify the factuality of the output responses and correct hallucinations. (Li et al. 2024b) automatically retrieve knowledge graphs to detect hallucinations through logical programming and mutation testing. Furthermore, (Bayat et al. 2023) propose an automated method to extract factual claims from responses and collect evidence from knowledge graphs to verify the factuality of the claims to be extracted. (2) Incorporating Alternative Models: Researchers use responses generated by other models for cross-validation to detect hallucinations (Hegselmann et al. 2024). (Cohen et al. 2023) construct a framework for assessing the factuality of output responses through cross-validation by two LMs. (Rawte et al. 2024; Wan et al. 2024) use multiple LLMs as ”judges” to evaluate various aspects of the model’s output responses. (Li et al. 2024b) propose a mutation testing model based on logic programming, which can verify the consistency of LLMs’ responses with real-world situations.
Black-box Hallucination Detection using Zero-resource
@@
Researchers propose using the model’s own capabilities to detect hallucinations (Liu et al. 2024b), because obtaining high-quality external resources is challenging (Mündler et al. 2023). (1) For non-long text responses generated by the model: Consistency comparison through multiple sampling responses is an important method (Allen, Polat, and Groth 2024). (Zhang et al. 2023) improve hallucination detection performance on commonsense QA tasks through semantic-aware cross-validation consistency. (Liu et al. 2024a) evaluate the reliability of responses generated by LLMs for individual questions or queries through cross-query comparison. (Ma, Dai, and Sui 2024) propose a critical calculation and conclusion (CCC) prompt template to enhance LLM’s ability to detect and correct unreasonable errors in mathematical problem solving. (Yehuda et al. 2024) identify potential instances of hallucination by quantifying the level of inconsistency between the original query and the reconstructed query. (2) For long text responses generated by the model: (Manakul, Liusie, and Gales 2023) propose a method to detect hallucinations by comparing the consistency of responses from multiple random samplings. (Yang, Sun, and Wan 2023) introduce a reverse validation method for passage-level hallucination detection in LLMs, which leverages the LLM’s own knowledge base. (Mündler et al. 2023) introduce a method for detecting self-contradictions in long text responses through logical reasoning. (Friel and Sanyal 2023) propose an efficient prompting method that uses the chains of thought generated by LLMs to evaluate and detect hallucinations in the long texts produced.
@@
Figure 1: GCA framework. We extract triples from the original response and sampled responses (left-upper concer). Then, we construct a graph for each response with the extracted triples and perform message passing and aggregation on the graph (as the upper branch). We conduct reverse validation for each part of the tripless with three reconstruction tasks (as the lower branch).
Method
Overview
@@
Our method consists of three modules: (1) Triple-Oriented Response Segmentation (Sec. 3.2) extracts knowledge facts from the model output responses. (2) Graph-based Contextual Consistency Comparison with RGCN (Sec. 3.3) constructs a graph carrying the extracted knowledge triples and utilizes an RGCN (relational graph convolutional network) (Schlichtkrull et al. 2018) to propagate and integrate messages across the graph. This method considers the dependencies between each knowledge triple (fact) and its surrounding triples during detection. (3) Reverse Verification via Triples Reconstruction (Sec. 3.4) achieves reverse verification for hallucination detection by reconstructing each triple via three LLM-based tasks.
@@
Triple-Oriented Response Segmentation
To better align knowledge facts in the consistency comparison, we propose to segment the responses by extracting knowledge triples as knowledge facts and checking the answers’ consistency among the triples. Our motivation is that due to the impact of wording, comparing textual consistency can lead to mismatches. Because hallucination detection considers the semantics of knowledge fact instead of specific word choices, we use a triple-based comparison method to provide better alignment than traditional textual comparison. Specifically, the steps are as follows:
@@
Graph-based Contextual Consistency Comparison with RGCN
To effectively consider dependencies between triples, we propose Graph-based Contextual Consistency Comparison (GCCC), which constructs a knowledge graph for each response and then conducts message passing and aggregation via RGCN. The intuition is that traditional consistency comparison focuses on comparing individual knowledge facts: it verifies a single piece of knowledge fact by comparing it only with the corresponding fact in the sampled responses at a time. It results in ignoring the triples that are mutually dependent on the given triple within the context information.
Therefore, we propose to construct graphs considering both the original response and the sampled responses. We employ RGCN for message passing and aggregation on these graphs. We divide the process into two stages: (1) knowledge triple modeling via graph learning. We build a graph for each response, and then we obtain node (entity) embeddings via RGCN processing to model the multiple knowledge triples for a response. (2) triples consistency comparison. We compare the consistency of triples across the graphs at the embedding level.
Knowledge Triple Modeling via Graph Learning

### v4

@@ Abstract
LLMs obtain remarkable performance but suffer from hallucinations. Most research on detecting hallucination focuses on questions with short and concrete correct answers that are easy to check faithfulness. Hallucination detections for text generation with open-ended answers are more hard. Some researchers use external knowledge to detect hallucinations in generated texts, but external resources for specific scenarios are hard to access. Recent studies on detecting hallucinations in long texts without external resources conduct consistency comparison among multiple sampled outputs. To handle long texts, researchers split long texts into multiple facts and individually compare the consistency of each pair of facts. However, these methods (1) hardly achieve alignment among multiple facts; (2) overlook dependencies between multiple contextual facts. In this paper, we propose a graph-based context-aware (GCA) hallucination detection method for text generations, which aligns facts and considers the dependencies between contextual facts in consistency comparison. Particularly, to align multiple facts, we conduct a triple-oriented response segmentation to extract multiple knowledge triples. To model dependencies among contextual triples (facts), we construct contextual triples into a graph and enhance triples’ interactions via message passing and aggregating via RGCN. To avoid the omission of knowledge triples in long texts, we conduct an LLM-based reverse verification by reconstructing the knowledge triples. Experiments show that our model enhances hallucination detection and excels all baselines [^1] Code:https://github.com/GCA-hallucinationdetection/GCA.
1 Introduction
@@
Recent research has shown that large language models (LLMs) achieved state-of-the-art performance in various NLP tasks (Fang et al. 2024; Lu et al. 2022). However, these models often suffer from hallucinations: generate incorrect or fabricated content in a factual way, which undermines models’ credibility (Ji et al. 2023; Lu et al. 2023a) and limits LLMs’ application in fields requiring factual accuracy (Huang et al. 2023; Su et al. 2024). Detecting hallucination in the model’s responses is crucial for LLMs’ boom.
@@
Existing hallucination detection studies mainly concentrate on tasks like question answering (QA) (Zhang et al. 2023a; Wen et al. 2023a) and arithmetic calculation (Xue et al. 2023; Tian et al. 2022) with short and concrete correct answers. In these tasks, the consistency among their concrete answers can be easily checked for hallucination (Jiang et al. 2021; Wang et al. 2024a). The research on detecting hallucinations in generated long text is more challenging because (1) generating text is open-ended and rarely has concrete answers and (2) long text contains multiple facts and the consistency among them is hard to verify. Therefore, we focus on hallucination detection in long text generation with black-box models (powerful LLM, like GPT-4) without external resources (i.e. zero-resource setting).
@@
Currently, the studies on black-box zero-resource hallucination detection for text generation can be divided into two categories: (1) Self-checking (Friel and Sanyal 2023; Liu et al. 2023) designs prompt texts using chain-of-thought (CoT) to verify response factuality by the LLMs’ own capabilities. Though that can be easily implemented and applied in various scenarios, it relies on the model’s own ability and can lead to missed detection: The model may overconfidently trust its outputs (Li et al. 2024a; Chen et al. 2024). (2) Consistency comparison (Zhang et al. 2023a; Ma et al. 2024) samples multiple responses to check whether the sampled responses are highly inconsistent, which indicates hallucination (Farquhar et al. 2024; Wang et al. 2024b). The method is effective for short responses with few concrete claims, making consistency comparison easy. However, in long text responses, diverse wording or lexical representations of the same fact complicate consistency comparison.
@@
To address those issues, researchers propose divide-and-conquer (Zhang et al. 2024) based on consistency comparison. It has three steps: (1) sample multiple additional responses appending to the original response; (2) divide each response into multiple facts; (3) compare the consistency of facts in the original response with those in the sampled responses, where highly inconsistent facts indicate hallucinations. However, it ignores the omission issue: facts in the original but not in sampled responses may be wrongly seen as hallucinations for lack of comparison. To overcome that, (Yang, Sun, and Wan 2023; Cao, Yang, and Zhao 2024) propose reverse verification. For each fact in the original response, they prompt LLMs to create a new question answered by the claim and check if the model’s response matches the fact. The method can avoid the omission issue because each fact has an accurately matched counterpart for comparison. However, a long text may contain multiple facts $\{A_{1},...A_{N}\}$, where $A_{i}$ is the $i$-th fact within LLMs’ generation. The above method compares the consistency of each fact sampled from different responses individually (e.g. verify the truthfulness of $A_{i}$ by comparing with $B_{i}$, where $B_{i}$ is $A_{i}$’s corresponding fact from a sampled response). It ignores the dependencies among each fact that can reduce detection performance. For example, for the response “Einstein won the Nobel Prize for proposing the theory of relativity.” we can extract two triples: (Einstein, proposal, theory of relativity) and (Einstein, receive, Nobel Prize). Although each triple is a fact, their dependency is an error. Therefore, considering the dependencies between a fact and its surrounding facts is promising to enhance hallucination detection.
@@
In this paper, we propose a graph-based context-aware (GCA) hallucination detection method for long text generation, which extracts triples to align facts for a better consistency comparison and considers the dependencies between contextual triples. Specifically, to better align facts from responses for consistency comparison, we conduct a triple-oriented response segmentation to extract multiple knowledge triples for each response. Triples are well-structured and thus easy for comparison; in addition, it is easy to align facts among multiple responses since triples carry only facts and get rid of the wording and lexical representations in the response (Wang et al. 2023b) (Sec. 3.2). To consider dependencies between a triple and its contextual triples, we construct a graph over a triple and its contextual triples. Then, we encourage triples’ interactions via message passing and aggregating features of neighboring nodes via RGCN (relational graph convolutional network) (Schlichtkrull et al. 2018) (Sec. 3.3). It encourages the interaction between facts as we observe that a certain number of facts has dependencies in a response (See Sec. 4.5). To avoid the omission issue in (Sec. 3.3), we propose an LLM-based reverse verification method with three reconstruction tasks to reconstruct the knowledge triples (Sec. 3.4). These tasks provide a more thorough and detailed detection of each triple in long text responses, enhancing the overall effectiveness of the method. Experiments show that our method effectively improves hallucination detection accuracy for long text responses generated by black-box models under zero-resource conditions.
@@
Our contributions are: (1) we propose a hallucination detection method for long text generation that considers the dependencies between contextual knowledge triples. (2) We propose a graph-based context-aware hallucination detection via consistency comparison with RGCN. (3) We additionally proposed three reversion verification tasks to help hallucination detection by reconstructing triples. (4) Experiments show that our method outperforms all baselines.
2 Related Work
2.1 White-box Hallucination Detection
@@
These methods analyze the model’s internal states to identify hallucinations (Yuksekgonul et al. 2024; Lu et al. 2023b; Wen et al. 2024), mainly divided into two types: (1) Output logit based method: The model’s output logits reflect the confidence of the model’s predictions (Jiang et al. 2024). (Varshney et al. 2023) calculates the logits for concepts in response and takes the minimal probabilities to model the uncertainty. (Verma et al. 2023) explores the integration of Epistemic Neural Networks (ENNs) with LLMs to improve the model’s uncertainty estimates. (Luo, Xiao, and Ma 2023) proposes to adjust the model’s output logits by adding a linear layer to better align with correctness likelihood. (2) Hidden layer activations based method: Hidden layer activations encapsulate the model’s internal representation of statement truthfulness (Fadeeva et al. 2024). (Azaria and Mitchell 2023) trains a classifier using LLM’s hidden layer activations to assess the truthfulness of statements. (Snyder, Moisescu, and Zafar 2024) uses output values from artifacts associated with model generation as input features to train a classifier that identifies hallucinations. (Zhu et al. 2024) uses probabilistic models to analyze internal state transitions in the LLM during generation to detect hallucinations.
2.2 Black-box Hallucination Detection using External Resources
@@
These methods aim to verify the authenticity of model-generated content by leveraging external knowledge (Wen et al. 2023b; Nahar et al. 2024). Depending on the source of external knowledge, it can be categorized into the following two types. (1) RAG-based method: Retrieval-augmented generation (RAG) is a technique that enhances text generation by retrieving relevant information from external sources (Sadat et al. 2023; Wang et al. 2023a). (Roychowdhury et al. 2024) proposes a multi-LLM system with the capability to retrieve external knowledge bases and perform real-time content authentication. (Ding et al. 2024) retrieves relevant evidence to help LLMs correct potential hallucinations in responses. (Kang, Ni, and Yao 2024) proposes a method for the real-time retrieval of Web search engines that can verify the factuality of the output responses and correct hallucinations. (Li et al. 2024b) automatically retrieves knowledge graphs to detect hallucinations through logical programming and mutation testing. Furthermore, (Bayat et al. 2023) proposes an automated method to extract factual claims from responses and collect evidence from knowledge graphs to verify the factuality of the claims to be extracted. (2) Incorporating Alternative Models: Researchers use responses generated by other models for cross-validation to detect hallucinations (Hegselmann et al. 2024). (Cohen et al. 2023) constructs a framework for assessing the factuality of output responses through cross-validation by two language models. (Rawte et al. 2024; Wan et al. 2024) use multiple LLMs as “judges” to evaluate various aspects of the model’s output responses. (Li et al. 2024b) proposes a mutation testing model based on logic programming, which can verify the consistency of LLMs’ responses with real-world situations.
2.3 Black-box Hallucination Detection using Zero-resource
@@
Researchers propose using the model’s own capabilities to detect hallucinations (Liu et al. 2024) because obtaining high-quality external resources is challenging (Mündler et al. 2024). (1) For non-long text responses generated by the model: Consistency comparison through multiple sampling responses is an important method (Allen, Polat, and Groth 2024). (Zhang et al. 2023a) improves hallucination detection performance on commonsense QA tasks through semantic-aware cross-validation consistency. (Liu et al. 2024) evaluates the reliability of responses generated by LLMs for individual questions or queries through cross-query comparison. (Ma et al. 2024) proposes a critical calculation and conclusion (CCC) prompt template to enhance LLM’s ability to detect and correct unreasonable errors in mathematical problem-solving. (Yehuda et al. 2024) identifies potential instances of hallucination by quantifying the level of inconsistency between the original query and the reconstructed query. (2) For long text responses generated by the model: (Manakul, Liusie, and Gales 2023) proposes a method to detect hallucinations by comparing the consistency of responses from multiple random samplings. (Yang, Sun, and Wan 2023) introduces a reverse validation method for passage-level hallucination detection in LLMs, which leverages the LLM’s own knowledge base. (Mündler et al. 2024) introduces a method for detecting self-contradictions in long text responses through logical reasoning. (Friel and Sanyal 2023) proposes an efficient prompting method that uses the chains of thought generated by LLMs to detect hallucinations in the responses. Unlike these LLM-based methods, our approach constructs long text responses as graphs and uses graph neural networks to capture the contextual influence of each fact during hallucination detection.
@@
Figure 1: GCA framework. We extract triples from the original response and sampled responses (left-upper corner). Then, we construct a graph for each response with the extracted triples and perform message passing and aggregation on the graph (as the upper branch). We conduct reverse validation for each part of the triples with three reconstruction tasks (as the lower branch).
3 Method
3.1 Overview
@@
Our method has three modules: (1) Triple-Oriented Response Segmentation (Sec. 3.2) extracts facts from the model’s responses. (2) Graph-based Contextual Consistency Comparison with RGCN (Sec. 3.3) constructs a graph carrying the extracted knowledge triples and utilizes an RGCN to propagate and integrate messages across the graph. It considers the dependencies between each knowledge triple (facts) and its surrounding triples during detection. (3) Reverse Verification via Triples Reconstruction (Sec. 3.4) achieves reverse verification for hallucination detection by reconstructing each triple via three LLM-based tasks (as shown in Fig.1).
@@
3.2 Triple-Oriented Response Segmentation
To better align facts in the consistency comparison, we propose to segment the responses by extracting knowledge triples as facts and checking the answers’ consistency among the triples. Our motivation is that due to the impact of wording, comparing textual consistency can lead to mismatches. Because hallucination detection considers the semantics of knowledge fact instead of specific word choices, we use a triple-based comparison method to provide better alignment than traditional textual comparison. Specifically, the steps are as follows:
@@ 3.2 Triple-Oriented Response Segmentation
3.3 Graph-based Contextual Consistency Comparison with RGCN
To effectively consider dependencies between triples, we propose Graph-based Contextual Consistency Comparison (GCCC), which constructs a knowledge graph for each response and then conducts message passing and aggregation via RGCN. The intuition is that traditional consistency comparison focuses on comparing individual facts: it verifies a single piece of knowledge fact by comparing it only with the corresponding fact in the sampled responses at a time. It results in ignoring the triples that are mutually dependent on the given triple within the context information.
To address the problem, our approach constructs graphs for the original response and the sampled responses. Then, it employs RGCN for message passing and aggregation on these graphs. The process consists of two stages: (1) knowledge triples modeling via graph learning. We build a graph for each response and obtain node (entity) embeddings via RGCN processing to model the multiple knowledge triples for a response. (2) triples consistency comparison. We compare the consistency of triples across the graphs at the embedding level.
Knowledge Triples Modeling via Graph Learning